OpenStudy (anonymous):

Find the points on the curve xy+4=0 at which the tangents are inclined at an angle of 45 with the x axis

OpenStudy (amistre64):

might want to start by taking a first derivative

OpenStudy (anonymous):

it comes y+4=0

OpenStudy (anonymous):

then?

OpenStudy (amistre64):

its not y+4 = 0

OpenStudy (anonymous):

oh..its the prod. rule pellet! back to you in a minute

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

its also good to know that slope = y/x, and tan(theta) = y/x are equivalent tan(45) = 1, so you are looking for where f'(c) = 1

OpenStudy (anonymous):

yea after these its coming dy/dx=-(y/x) what do i do now?

OpenStudy (anonymous):

ok done :D thanks mate :D :D

OpenStudy (anonymous):

really thanks

OpenStudy (amistre64):

youre welcome